OpenStudy (anonymous):

Can you check if i rationalized the denominator correctly? (its in the comments)

OpenStudy (anonymous):

\[-3/\sqrt{x}-4\] is \[-3\sqrt{x}+4/x-16\]

OpenStudy (anonymous):

This is not correct form: \[= \frac{-3 \times \sqrt{x}}{\sqrt{x} \times \sqrt{x}} - 4 \implies \frac{-3\sqrt{x}}{x} - 4\]

OpenStudy (anonymous):

\[-3 \over \sqrt{x}-4\] is \[3\sqrt{x}+4 \over x-16\]. Better?

OpenStudy (anonymous):

\[= \frac{-3 \times (\sqrt{x}+4)}{(\sqrt{x}-4) \times (\sqrt{x} +4)} - 4 \implies \frac{-3\sqrt{x} -12}{x-16}\] I got this much..

OpenStudy (anonymous):

I got -12 first too but i checked it with mathway so now i have no idea which one is correct.

OpenStudy (anonymous):

The above that has written by me is correct.. You can trust me..

OpenStudy (anonymous):

OK, if you say so.

OpenStudy (anonymous):

I am not saying dear.. I have done it very correctly for you..

OpenStudy (phi):

When you type \[ -3/\sqrt{x}-4 \] that means \[ \frac{-3}{\sqrt{4}}-4 \] on the other hand (with parentheses) \[ -3/(\sqrt{x}-4)= \frac{-3}{\sqrt{x}-4} \] So the answer depends on which is the real question.

OpenStudy (anonymous):

I rewrote it in the fourth post.